# Diversity-Enhanced Reasoning for Subjective Questions

### v1

@@ 8 Conclusion
We introduce MultiRole-R1, a novel framework that enhances the reasoning capabilities of Large Reasoning Models (LRM) on subjective tasks. Our framework enables models to think from multiple perspectives before providing a final answer, enabling diverse answers for subjective questions. To achieve this, we employ Group Relative Policy Optimization (GRPO) with a shaped reward that combines a verifiable accuracy signal with a diversity metric. Furthermore, through training on subjective questions solely, we are able to enhance the models’ performance on both subjective and objective reasoning. Our analysis reveals that diversity is a more consistent indicator of performance than response length, showing a strong per-task correlation with accuracy that significantly outweighs the correlation with length. This suggests that the improvements in modern reasoning models like o1 or R1 are driven by a scaling law of diversity, not just superficial verbosity. Our work underscores the importance of diversity at both perspective and lexical levels, paving new directions for future research on diversity-enhance training of LRMs.
Limitations
One of the main limitations of our study is that due to the constraints of computational resources, we are unable to scale to larger models, e.g. models of 70B. For the multirole supervised finetuning process, since the filtering results after self-consistency are still noisy, the final alignment performance may be constrained by the noisy labels in the positive samples. Our role generation solely depends on the model itself, which may not accurately represent the opinion of the real social community. Future work can further investigate the model’s culture and opinion alignment to different demographics, by incorporating the value of a broader demographic of users. In addition, beyond text-only data, extending diversity-enhanced reasoning in the multimodal domain through thinking with images Su et al. (2025) stands as another underexplored yet meaningful research area to investigate further.
Ethics Statements
Our study focuses on subjective question answering, which encourages LLM’s answers to include more diverse social identities, moral values, and nationalities. To achieve multiple-perspective reasoning, our approach utilizes the model’s own outputs to perform role generation. We also unleash the LLM’s role-playing ability to generate diverse opinions based on the role perspectives. By reducing dependence on manual labeling, this method enhances the fairness, scalability, and inclusivity of AI, which will further the democratization of LLMs across broader demographics.
Author Contribution Statements
@@
Appendix A Training and Testing Data Split
@@
Table 6: Statistic of the number of training data after self-consistency filtering (consistency filter) and ground-truth-guided hinted filtering (GT filter). We also report the number of test data used in the evaluation phase.
Appendix B Qualitative Examples
B.1 Qualitative Examples of Test Time Scaling Set Up
Table 7 shows the different lengths of the reasoning chain in the pilot analysis. Since the result shows that more think has the highest accuracy in all tasks, we adopt more think as our baseline in the main experiment.
@@ B.1 Qualitative Examples of Test Time Scaling Set Up
Output 111 Okay, so the question is whether online courses are more effective than traditional classrooms.
@@
Output 222 Let me take teaching effectiveness, student engagement, workload, and feedback quality into consideration.
@@
Output $N$ Employers in tech may value online certifications.
@@
B.2 Qualitative Example of Different Roles
@@
Figure 5: We present the top 100 most frequent roles from our SFT training dataset out of the total 968 roles, which are generated from the questions in the training data. The diameters of the circles are proportional to the frequency.
Appendix C Prompts
C.1 Role Generation Prompt
@@
C.2 Evaluation Prompts
@@
C.3 Role Generation Prompts
We show the prompts for role generation in Figure 6.
@@ C.3 Role Generation Prompts
Appendix D SFT Training Configuration
@@
Table 8: SFT training parameter and inference parameter.
Appendix E Diversity Metric
E.1 Formulation of the Diversity Metric
@@
  • Lexical Diversity ($D_{lex}$): Measures the variety of unique words in a text using the Type-Token Ratio (TTR) Hess et al. (1984), reflecting vocabulary richness.
  • Entropy Diversity ($D_{ent}$): Evaluates the unpredictability of word usage based on information entropy Kalimeri et al. (2014), capturing distributional variety.
  • Sentence Length Diversity ($D_{len}$): Assesses variation in sentence lengths using the coefficient of variation Chen et al. (2020), indicating structural diversity.
  • Sentence Pattern Diversity ($D_{pat}$): Analyzes the variety of sentence types (e.g., declarative, interrogative) and their distribution Shaib et al. (2025).
  • Adjacent Sentence Diversity ($D_{yule}$): Examines differences between consecutive sentences using Jaccard distance Farouk (2019), highlighting contextual shifts.
  • Yule’s K Diversity ($D_{bi}$): Reflects vocabulary distribution by analyzing word frequency patterns Tanaka-Ishii and Aihara (2015), with lower values indicating higher diversity.
  • Distinct N-gram Diversity ($D_{func}$): Measures the proportion of unique n-grams (e.g., bigrams) in the text Li et al. (2016), showcasing phrase-level variety.
  • Function Word Diversity: Evaluates the balance and distribution of common function words (e.g., articles, prepositions) to assess linguistic variety Segarra et al. (2015).

### v3

@@ 7 Conclusion
We introduce MultiRole-R1, a diversity-enhanced training framework that enhances the reasoning capabilities of Large Reasoning Models (LRM) by optimizing perspective diversity and token-level diversity, by self-synthesizing multi-role reasoning paths and incorporating diversity reward shaping. By training exclusively on subjective questions, MultiRole-R1 demonstrates robust generalization to out-of-domain subjective and objective tasks, including advanced mathematics. By taking token-level diversity as reward shaping, we broaden the search space of Chain-of-Thought (CoT), as evidenced by an increase in pass@k accuracy. Our analysis investigates the validity of our design choices, revealing that diversity is a more reliable indicator of accuracy than superficial verbosity, also demonstrating that MultiRole-R1 enables efficient reasoning. Our findings highlight that prioritizing diversity in (CoT) is more effective than simply lengthening the reasoning chain, showing promising directions for subjective reasoning enhancement.
Large Language Model Usage Statement
In this work, a Large Language Model (LLM) was utilized as a writing assistant. The authors provided their draft to the LLM for suggestions to improve grammar, enhance phrasing clarity, and remove non-academic language. The model was also used to brainstorm potential titles. The final manuscript, including the title, was determined and refined by the authors, who retained full editorial control.
Author Contributions
@@
Appendix A Theoretical Derivations
We provide step-by-step derivations for core components of our framework as follows:
A.1 Entropy-Regularized Role Selection Yields Softmax
Let $S_{i}(\mathcal{Q})\triangleq\mathbb{E}[\mathcal{M}(\mathcal{R}_{i}|\mathcal{Q})]+\alpha\mathbb{E}_{j\neq i}[1-\mathrm{sim}(\mathcal{R}_{i},\mathcal{R}_{j})]$ be the relevance-plus-contrast score for role $\mathcal{R}_{i}$, and let $\mathbf{p}\in\Delta_{n}$ be a probability vector over roles. Consider the entropy-regularized objective with temperature $\eta>0$:
$\max_{\mathbf{p}\in\Delta_{n}}\sum_{i=1}^{n}p_{i}S_{i}(\mathcal{Q})+\eta H(\mathbf{p}),\quad\text{where}\quad H(\mathbf{p})=-\sum_{i=1}^{n}p_{i}\log p_{i}.$ (5)
Form the Lagrangian with multiplier $\lambda$ for $\sum_{i}p_{i}=1$:
$\mathcal{L}(\mathbf{p},\lambda)$ $=\sum_{i=1}^{n}p_{i}S_{i}(\mathcal{Q})-\eta\sum_{i=1}^{n}p_{i}\log p_{i}+\lambda\left(\sum_{i=1}^{n}p_{i}-1\right)$ (6)
$\frac{\partial\mathcal{L}}{\partial p_{i}}$ $=S_{i}(\mathcal{Q})-\eta(1+\log p_{i})+\lambda\stackrel{{\scriptstyle!}}{{=}}0$ (7)
$\Rightarrow\log p_{i}$ $=\frac{S_{i}(\mathcal{Q})+\lambda-\eta}{\eta}\quad\Longrightarrow\quad p_{i}=\exp\left(\frac{S_{i}(\mathcal{Q})}{\eta}\right)\cdot\exp\left(\frac{\lambda-\eta}{\eta}\right).$ (8)
Impose normalization $\sum_{i}p_{i}=1$:
$\sum_{i=1}^{n}p_{i}=\exp\left(\frac{\lambda-\eta}{\eta}\right)\sum_{i=1}^{n}\exp\left(\frac{S_{i}(\mathcal{Q})}{\eta}\right)=1,$ (9)
which yields
$\exp\left(\frac{\lambda-\eta}{\eta}\right)=\frac{1}{\sum_{k=1}^{n}\exp\big(S_{k}(\mathcal{Q})/\eta\big)}.$ (10)
Substituting back,
$p_{i}^{*}=\frac{\exp\big(S_{i}(\mathcal{Q})/\eta\big)}{\sum_{k=1}^{n}\exp\big(S_{k}(\mathcal{Q})/\eta\big)}.$ (11)
Setting $\eta=1$ recovers the softmax policy $P(\mathcal{R}_{i}|\mathcal{Q})=\mathrm{softmax}\big(S_{i}(\mathcal{Q})\big)$ used in our method.
A.2 Self-Consistency Filtering as MAP under Dirichlet–Multinomial
For role $\mathcal{R}_{i}$, let $\{\mathcal{T}_{\mathcal{R}_{i}}^{(j)}\}_{j=1}^{k}$ be $k$ samples grouped into $K$ semantic equivalence classes $\{\mathcal{C}_{\ell}\}_{\ell=1}^{K}$ with counts $n_{\ell}$. Assume a symmetric Dirichlet prior $\bm{\theta}\sim\mathrm{Dir}(\alpha,\dots,\alpha)$ over class probabilities and multinomial likelihood. The posterior is
$p(\bm{\theta}\mid\{n_{\ell}\})=\mathrm{Dir}\big(\alpha+n_{1},\dots,\alpha+n_{K}\big).$ (12)
The mode (MAP estimate) of a Dirichlet with parameters $\beta_{\ell}=\alpha+n_{\ell}$ is
$\theta_{\ell}^{\mathrm{MAP}}=\frac{\beta_{\ell}-1}{\sum_{m=1}^{K}\beta_{m}-K}=\frac{\alpha+n_{\ell}-1}{K\alpha+\sum_{m=1}^{K}n_{m}-K}.$ (13)
Thus, the class with largest MAP component is
$\hat{\ell}_{\mathrm{MAP}}=\arg\max_{\ell}\theta_{\ell}^{\mathrm{MAP}}=\arg\max_{\ell}\big(\alpha+n_{\ell}-1\big).$ (14)
For $\alpha\geq 1$, the mapping $\ell\mapsto\alpha+n_{\ell}-1$ is monotonically increasing in $n_{\ell}$, hence
$\hat{\ell}_{\mathrm{MAP}}=\arg\max_{\ell}n_{\ell},$ (15)
which is precisely the majority-vote rule. Therefore, our self-consistency filter returns the MAP-equivalent class under a symmetric Dirichlet prior.
For binary equivalence ($K=2$) with success probability $\theta>1/2$, the probability that majority vote errs satisfies Hoeffding’s inequality:
$\Pr\left[\sum_{j=1}^{k}Y_{j}\leq\frac{k}{2}\right]$ $=\Pr\left[\frac{1}{k}\sum_{j=1}^{k}Y_{j}-\theta\leq\frac{1}{2}-\theta\right]$ (16)
$\leq\exp\Big(-2k(\theta-1/2)^{2}\Big),$ (17)
where $Y_{j}=\mathbbm{1}_{\mathcal{T}_{\mathcal{R}_{i}}^{(j)}\in\mathcal{C}_{\text{canonical}}}$. Hence, consistency filtering is exponentially reliable in $k$ under mild conditions.
A.3 Potential-Based Diversity Shaping Preserves Optimal Policies
Let the shaped reward be $\mathbf{R}=\delta\mathbf{R}_{\text{acc}}+(1-\delta)\mathbf{R}_{\text{div}}$ with $\delta\in(0,1)$, and define a potential $\Phi:\mathcal{S}\to\mathbb{R}$ over states (prefixes of multi-role traces) such that
$\mathbf{R}_{\text{div}}(s_{t},a_{t},s_{t+1})\triangleq\gamma\Phi(s_{t+1})-\Phi(s_{t}),\quad 0<\gamma<1,\quad\Phi(s_{T})=0.$ (18)
For any trajectory $(s_{0},a_{0},\dots,s_{T})$, the discounted return is
$\sum_{t=0}^{T-1}\gamma^{t}\mathbf{R}(s_{t},a_{t},s_{t+1})$ (19)
$=\delta\sum_{t=0}^{T-1}\gamma^{t}\mathbf{R}_{\text{acc}}(s_{t},a_{t},s_{t+1})+(1-\delta)\sum_{t=0}^{T-1}\gamma^{t}\big(\gamma\Phi(s_{t+1})-\Phi(s_{t})\big)$ (20)
$=\delta\sum_{t=0}^{T-1}\gamma^{t}\mathbf{R}_{\text{acc}}(s_{t},a_{t},s_{t+1})+(1-\delta)\left(\sum_{t=0}^{T-1}\gamma^{t+1}\Phi(s_{t+1})-\sum_{t=0}^{T-1}\gamma^{t}\Phi(s_{t})\right)$ (21)
$=\delta\sum_{t=0}^{T-1}\gamma^{t}\mathbf{R}_{\text{acc}}(s_{t},a_{t},s_{t+1})+(1-\delta)\left(\gamma^{T}\Phi(s_{T})-\Phi(s_{0})\right)$ (22)
$=\delta\sum_{t=0}^{T-1}\gamma^{t}\mathbf{R}_{\text{acc}}(s_{t},a_{t},s_{t+1})-(1-\delta)\Phi(s_{0}),$ (23)
which differs from the accuracy-only return by the constant $(1-\delta)\Phi(s_{0})$ that is independent of actions. Therefore, diversity shaping via potentials preserves the set of optimal policies.
A.4 GRPO Advantages: Degeneracy and Variance Recovery from Diversity
In GRPO, for a group of $G$ samples $\{y_{i}\}_{i=1}^{G}$ with rewards $\{\mathbf{R}_{i}\}_{i=1}^{G}$, define
$\mu=\frac{1}{G}\sum_{i=1}^{G}\mathbf{R}_{i},\quad\sigma^{2}=\frac{1}{G}\sum_{i=1}^{G}(\mathbf{R}_{i}-\mu)^{2},\quad\mathbf{A}_{i}=\frac{\mathbf{R}_{i}-\mu}{\sigma}.$ (24)
The surrogate loss and gradient are
$\mathcal{L}(\pi)=-\frac{1}{G}\sum_{i=1}^{G}\mathbf{A}_{i}\log\pi(y_{i}\mid x),\quad\nabla\mathcal{L}(\pi)=-\frac{1}{G}\sum_{i=1}^{G}\mathbf{A}_{i}\nabla\log\pi(y_{i}\mid x).$ (25)
If all rewards are equal, $\mathbf{R}_{i}=c$, then
$\mu=c,\quad\sigma^{2}=\frac{1}{G}\sum_{i}(c-c)^{2}=0\quad\Longrightarrow\quad\mathbf{A}_{i}=0\quad\Longrightarrow\quad\nabla\mathcal{L}(\pi)=\mathbf{0},$ (26)
which stalls learning.
With shaped rewards $\mathbf{R}_{i}=\delta\mathbf{R}_{\text{acc},i}+(1-\delta)\mathbf{R}_{\text{div},i}$, the group variance expands as
$\mathrm{Var}[\mathbf{R}]$ $=\mathrm{Var}\big[\delta\mathbf{R}_{\text{acc}}+(1-\delta)\mathbf{R}_{\text{div}}\big]$ (27)
$=\delta^{2}\mathrm{Var}[\mathbf{R}_{\text{acc}}]+(1-\delta)^{2}\mathrm{Var}[\mathbf{R}_{\text{div}}]+2\delta(1-\delta)\mathrm{Cov}(\mathbf{R}_{\text{acc}},\mathbf{R}_{\text{div}}).$ (28)
In the extreme case $\mathrm{Var}[\mathbf{R}_{\text{acc}}]=0$ (all-correct or all-incorrect within the group),
$\mathrm{Var}[\mathbf{R}]=(1-\delta)^{2}\mathrm{Var}[\mathbf{R}_{\text{div}}]+2\delta(1-\delta)\mathrm{Cov}(\mathbf{R}_{\text{acc}},\mathbf{R}_{\text{div}}).$ (29)
If $\mathrm{Var}[\mathbf{R}_{\text{div}}]>0$ and the covariance is not exactly $\frac{1-\delta}{\delta}\mathrm{Var}[\mathbf{R}_{\text{div}}]$, then $\mathrm{Var}[\mathbf{R}]>0$, guaranteeing nonzero advantages and informative gradients.
A.5 Diversity Reduces Correlation and Tightens Ensemble Accuracy Bounds
Let $Z_{i}\in\{0,1\}$ indicate correctness of role $i$, with $\Pr[Z_{i}=1]=p>1/2$ and pairwise correlation $\rho=\mathrm{Corr}(Z_{i},Z_{j})$ (exchangeable roles). Define $S_{m}=\sum_{i=1}^{m}Z_{i}$ and majority vote $\hat{Z}=\mathbbm{1}_{S_{m}>m/2}$. The mean and variance of $S_{m}$ are
$\mu=\mathbb{E}[S_{m}]=mp,\quad\sigma^{2}=\mathrm{Var}[S_{m}]=\sum_{i=1}^{m}\mathrm{Var}[Z_{i}]+2\sum_{1\leq i<j\leq m}\mathrm{Cov}(Z_{i},Z_{j}).$ (30)
Using $\mathrm{Var}(Z_{i})=p(1-p)$ and $\mathrm{Cov}(Z_{i},Z_{j})=\rho\cdot p(1-p)$,
$\sigma^{2}$ $=m\cdot p(1-p)+2\cdot\frac{m(m-1)}{2}\cdot\rho\cdot p(1-p)$ (31)
$=p(1-p)\big(m+m(m-1)\rho\big).$ (32)
We bound the one-sided tail via Cantelli’s inequality with $t=\mu-m/2=m(p-1/2)>0$:
$\Pr\left[S_{m}\leq\frac{m}{2}\right]$ $=\Pr\big[S_{m}-\mu\leq-t\big]$ (33)
$\leq\frac{\sigma^{2}}{\sigma^{2}+t^{2}}$ (34)
$=\frac{p(1-p)\big(m+m(m-1)\rho\big)}{p(1-p)\big(m+m(m-1)\rho\big)+m^{2}(p-1/2)^{2}}.$ (35)
The right-hand side is monotone increasing in $\rho$ because it is an increasing function of $\sigma^{2}$. Therefore, reducing positive correlation (increasing diversity) tightens the bound on ensemble error, improving majority-vote accuracy.
Appendix B Training Set Up
B.1 Training and Testing Data Split
@@
Table 6: Statistics of the number of training data after self-consistency filtering (consistency filter) and ground-truth-guided hinted filtering (GT filter). We also report the number of test data used in the evaluation phase.
@@
B.2 Hyperparameter Setting
In this section, we discuss the choice of diversity reward shaping hyperparameter $\gamma$. We pilot the GRPO training the DeepSeek-R1-Distill-Qwen-7B model. Among common empirical $\gamma$ value of $\gamma=0.05,0.1,0.2,0.3$, we found that $\gamma=0.1$ yields the highest accuracy of 49.1% in Global Opinion QA. This is also consistent with Ng et al. (1999), and we applied the same parameter for other models and settings.
Appendix C Qualitative Examples
C.1 Qualitative Examples of Subjective Questions Ground Truth Format
Since the subjective questions have multiple valid answers based on different roles, we provide a concrete example of the ground-truth format, taking the Global Opinion QA dataset as an example, if if there are choices A, B, C and D, the ground truth look like the following: {‘Belgium’: {‘A’: 0.21, ‘B’: 0.07, ‘C’: 0.69, ‘D’: 0.03}, ‘France’: {‘A’: 0.54, ‘B’: 0.09, ‘C’: 0.35, ‘D’: 0.02}, ...}. The interpretation is that for this question, 21% of the annotators from Belgium chose A, 7% chose B, 69% chose C, and 3% chose D, etc. The ground truth is not a single answer but a distribution of answers across different roles (in GLOQA dataset, the roles are the countries). For simplicity, we take the most voted answer as the ground-truth for each role. In the above example, ‘C’ is the ground-truth of Belgium.
C.2 Qualitative Examples of Test Time Scaling Set Up
Table 7 shows the different lengths of the reasoning chain in the pilot analysis. Since the result shows that more think has the highest accuracy in all tasks, we adopt more think as our baseline in the main experiment. The definition of different reasoning set-up is as follows:
@@ C.2 Qualitative Examples of Test Time Scaling Set Up
Output 11 Okay, so the question is whether online courses are more effective than traditional classrooms.
@@
Output 22 Let me take teaching effectiveness, student engagement, workload, and feedback quality into consideration.
@@
Output NN Employers in tech may value online certifications.
@@
C.3 Qualitative Example of Different Roles
@@
Figure 5: We present the top 100 most frequent roles from the model output during test-time. The diameters of the circles are proportional to the frequency.
Appendix D Prompts
D.1 Role Generation Prompt
@@
D.2 Role Generation Prompts
We show the prompts for role generation in Figure 6.
D.3 Evaluation Prompts
@@
Appendix E SFT Training Configuration
@@
Appendix F Diversity Metric
F.1 Formulation of the Diversity Metric
@@
  • Lexical Diversity ($D_{\mathrm{lex}}$): Type–Token Ratio (TTR) (Hess et al., 1984). Let the text have $N$ tokens and vocabulary $V$ of unique types.
    $D_{\mathrm{lex}}\;=\;\frac{|V|}{N}.$
  • Entropy Diversity ($D_{\mathrm{ent}}$): Normalized token entropy (Kalimeri et al., 2014). Let $p(w)$ be the empirical probability of word $w\in V$.
    $D_{\mathrm{ent}}\;=\;-\frac{1}{\log|V|}\sum_{w\in V}p(w)\,\log p(w).$
  • Sentence Length Diversity ($D_{\mathrm{len}}$): Coefficient of variation of sentence lengths (Chen et al., 2020). For $S$ sentences with lengths (in tokens) $\ell_{1},\dots,\ell_{S}$, mean $\bar{\ell}=\tfrac{1}{S}\sum_{i}\ell_{i}$, and standard deviation $\sigma_{\ell}=\sqrt{\tfrac{1}{S}\sum_{i}(\ell_{i}-\bar{\ell})^{2}}$,
    $D_{\mathrm{len}}\;=\;\frac{\sigma_{\ell}}{\bar{\ell}}.$
  • Sentence Pattern Diversity ($D_{\mathrm{pat}}$): Normalized entropy over sentence types (e.g., declarative, interrogative) (Shaib et al., 2025). With type set $\mathcal{C}$ and proportions $p_{c}$ over $c\in\mathcal{C}$,
    $D_{\mathrm{pat}}\;=\;-\frac{1}{\log|\mathcal{C}|}\sum_{c\in\mathcal{C}}p_{c}\,\log p_{c}.$
  • Adjacent Sentence Diversity ($D_{\mathrm{adj}}$): Mean Jaccard distance between consecutive sentences (Farouk, 2019). Let $A_{i}$ be the (lemmatized or lowercased) token set of sentence $i$.
    $D_{\mathrm{adj}}\;=\;\frac{1}{S-1}\sum_{i=1}^{S-1}\Bigl(1-\frac{|A_{i}\cap A_{i+1}|}{|A_{i}\cup A_{i+1}|}\Bigr).$
  • Yule’s K ($K_{\mathrm{Yule}}$): Vocabulary concentration based on frequency spectrum (Tanaka-Ishii & Aihara, 2015). Let $V_{r}$ be the number of types that occur exactly $r$ times and $N$ the total tokens.
    $K_{\mathrm{Yule}}\;=\;10^{4}\,\frac{\sum_{r\geq 1}r^{2}V_{r}-N}{N^{2}}\quad\text{(lower $\Rightarrow$ higher diversity)}.$
    (Optional normalized variant: $D_{\mathrm{yule}}=\frac{1}{1+K_{\mathrm{Yule}}}$.)
  • Distinct $n$-gram Diversity ($D_{\mathrm{dist}}^{(n)}$): Proportion of unique $n$-grams (Li et al., 2016). If the text has $M_{n}$ total $n$-grams and $U_{n}$ unique $n$-grams,
    $D_{\mathrm{dist}}^{(n)}\;=\;\frac{U_{n}}{M_{n}}\qquad\text{(optionally average over $n\in\{1,2,3\}$)}.$
  • Function Word Diversity ($D_{\mathrm{func}}$): Normalized entropy over a fixed function-word inventory $\mathcal{F}$ (articles, prepositions, etc.) (Segarra et al., 2015). With counts $f_{j}$ and probabilities $p_{j}=\tfrac{f_{j}}{\sum_{k\in\mathcal{F}}f_{k}}$ for $j\in\mathcal{F}$,
    $D_{\mathrm{func}}\;=\;-\frac{1}{\log|\mathcal{F}|}\sum_{j\in\mathcal{F}}p_{j}\,\log p_{j}.$
F.2 Embedding Similarity as the Diversity Metric
$RODS=\frac{2}{n(n-1)}\sum_{i<j}d_{ij}$
Table 9: Detailed composition of the diversity scores based on the output of R1-Distilled-Qwen-7B. This includes lexical, entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, bigram, and the function word diversity score across all tasks and baseline settings. In addition, we also provide the combined diversity score, average reasoning length and length normalized diversity score.